# A meshless method for computational electromagnetics with improved dispersion properties

### v1

@@ 7 Conclusion
The starting point of our research was a seemingly simple question - How can the FDTD method be generalised to a scattered setting and would this help with the dispersion anisotropy? The first question was answered affirmatively with the use of RBF-based meshless methods - we have performed a straightforward generalisation of the FDTD by simply replacing the derivative operators appropriately. We have considered two approaches, which a-priori both seemed equally promising: The usual RBF-FD and the virtual stencil approach, in this paper referred to as RBF-VFD. Unlike FDTD, which enjoys a simple stability criterion, major issues were found regarding stability of the proposed methods and we have demonstrated that these issues can be overcome by employing hyperviscosity stabilisation, describing a simple framework on how the appropriate HV parameters should be selected.
We then analysed the dispersion properties and found that RBF-FD possesses a parameter that can control the amount of dispersion - the stencil size. Moving to the angular dependence of the dispersion, we have answered affirmatively also to the second part of our question - meshless methods can eliminate dispersion anisotropy, which in our example happened with the RBF-FD method. However, this is not always the case, as demonstrated by the RBF-VFD approach.
@@
Several important points were glossed over in our research that should be addressed if the proposed methods are ever to be used in an industrial setting. First, our analysis was limited to an effectively one-dimensional case. As part of our current research, we are applying those same methods to tackle different cases, such as scattering on an irregular body or some 3D problems. This also includes analyses of the behaviour of the method, when applied to cases with spatially varying discretisation density.
An important advantage of the FDTD method that was not given much attention is its intrinsic divergence conservation [12]. In this aspect, the RBF-VFD, being in some sense closer to the FDTD, could perform better compared to the RBF-FD approach, especially if paired with the so-called divergence-free RBFs [17], paving the way for potential future work in that direction.
Finally, our observations of the stencil dependence of dispersion remain unexplained. However, we strongly suspect that the reason behind it can be connected to the recent developments in the RBF-FD method [31], where it was shown that the basis of our interpolant in the global sense is discontinuous, with the jumps decreasing with the stencil size. Further work on this matter could put our observations in a concrete, mathematical setting.
@@ Acknowledgements
Furthermore, the authors would like to thank the following individuals, listed in alphabetical order, for their assistance with various aspect of this study: Vito Levstik, Vagif Osmanov, Adrijan Rogan, Miha Rot and Žiga Vaupotič.
8 Appendix: Naive instability on scattered nodes
Here we aim to provide some intuition on why a naive approach to solving Maxwell’s equations on scattered nodes without any stabilisation fails.
@@ 8 Appendix: Naive instability on scattered nodes
$H_{x}(x_{i},y_{i},n+1)=H_{x}(x_{i},y_{i},n)-\frac{\Delta t}{\mu_{0}}D^{\partial_{y}}(E_{z}(x_{i},y_{i},n)),$ (27)
$H_{y}(x_{i},y_{i},n+1)=H_{y}(x_{i},y_{i},n)+\frac{\Delta t}{\mu_{0}}D^{\partial_{x}}(E_{z}(x_{i},y_{i},n)),$
@@
By Von Neumann we take an ansatz888a general Fourier component has $i(k_{x}x+k_{y}y)$ in the exponential. However, we will see that unstable modes arise already in this special case.:
@@
Consdering the update equations (27) for a single point $(x_{i},y_{i})$, plugging in the ansatz and dividing by $e^{ik(x+y)}$ we get:
@@
where we have introduced the notation:
@@
For stability, it is necessary that the spectral radius of the matrix in the last equation is below unity. We can calculate the eigenvalues to equal:
@@
We can expect unstability if one of the $\lambda_{2,3}$ has modulus greater than unity. Denoting $\xi\mathrel{\mathop{:}}=c_{1}c_{2}(w_{x}^{2}+w_{y}^{2})$ we can write:
@@
Observe that the product of the two equals unity:
$\lambda_{2}*\lambda_{3}=(1+0.5\xi)^{2}-0.5^{2}(\xi^{2}+4\xi)=1$ (35)
Therefore, $\lambda_{2}=1/\lambda_{3}$. From this it follows that if $|\lambda_{2}|<1$ then $|\lambda_{3}|>1$ and an unstable mode has been found. We must only exclude the possiblity of $|\lambda_{2}|=|\lambda_{3}|=1$. However, in this case we can write $\lambda_{2}=e^{i\alpha},\lambda_{3}=e^{-i\alpha}$ and calculate:
@@
As emphasised earlier, this is by no means a formal proof of the instability of any meshless approach, but instead merely an intuition as to where the locally unstable modes may come from. The analysis was limited to an interleaved leapfrog time-stepping, with no field staggering and a general expression of a derivative in a stencil. This includes our RBF-FD and RBF-VFD methods.

### v2

@@ 7 Conclusion
The starting point of our research was a seemingly simple question – how can the FDTD method be generalised to a scattered setting and would this help with the numerical dispersion anisotropy? The first question was answered affirmatively with the use of RBF-based meshless methods – we have performed a straightforward generalisation of the FDTD by simply replacing the derivative operators appropriately. We have considered two approaches, which a-priori both seemed equally promising: The usual RBF-FD and the virtual stencil approach, in this paper referred to as the RBF-VFD. Unlike FDTD, which enjoys a simple stability criterion, major issues were found regarding stability of the proposed methods and we have demonstrated that these issues can be overcome by employing hyperviscosity stabilisation, describing a simple framework on how the appropriate HV parameters should be selected.
We then analysed the spectral picture of the solution and found that our proposed methods possess a parameter that can control the amount of dispersion - stencil size. Moving to the angular dependence of the dispersion, we have answered affirmatively also to the second part of our initial question - meshless methods can eliminate dispersion anisotropy, which in our example happened with the RBF-FD method. However, this is not always the case, as demonstrated by the RBF-VFD approach.
@@
Several important points were glossed over in our research that should be addressed if the proposed methods are ever to be used in an industrial setting. First, our analysis was limited to an effectively one-dimensional case. As part of our current research, we are applying those same methods to tackle different cases, such as scattering on an irregular body or some 3D problems. This also includes analyses of the behaviour of the method when applied to cases with spatially varying discretisation density.
An important advantage of the FDTD method that was not given much attention is its intrinsic divergence conservation [12]. In this aspect, the RBF-VFD, being in some sense closer to the FDTD, could perform better compared to the RBF-FD approach, especially if paired with divergence-free RBFs [17], paving the way for potential future work in that direction.
Finally, the stencil dependence of numerical dispersion remains unexplained. We strongly suspect that the reason behind it can be connected to the recent developments in the RBF-FD method [32], where it was shown that the basis of our interpolant in the global sense is discontinuous, with the jumps decreasing with the stencil size. Further work on this matter could put our observations in a formal setting with supporting mathematical justifications.
@@ Acknowledgements
Furthermore, the authors would like to thank the following individuals for their assistance with various aspect of this study, in alphabetical order: Levstik Vito, Osmanov Vagif, Rogan Adrijan, Rot Miha and Vaupotič Žiga.
Appendix A Naive instability on scattered nodes
The purpose of this Appendix is to provide some intuition on why a naive approach to solving Maxwell’s equations on scattered nodes without any stabilisation fails.
@@ Appendix A Naive instability on scattered nodes
$H_{x}(x_{i},y_{i},n+1)=H_{x}(x_{i},y_{i},n)-\frac{\Delta t}{\mu_{0}}D^{\partial_{y}}(E_{z}(x_{i},y_{i},n)),$
$H_{y}(x_{i},y_{i},n+1)=H_{y}(x_{i},y_{i},n)+\frac{\Delta t}{\mu_{0}}D^{\partial_{x}}(E_{z}(x_{i},y_{i},n)),$ (28)
@@
We take an ansatz999a general Fourier component has $i(k_{x}x+k_{y}y)$ in the exponential. However, we will see that unstable modes arise already in this special case.:
@@
Considering the update equations (28) for a single point $(x_{i},y_{i})$, plugging in the ansatz and dividing by $e^{ik(x+y)}$ we get:
@@
where we have introduced:
@@
For stability, it is necessary that the spectral radius of the matrix in the last equation does not exceed unity. We can calculate the eigenvalues to equal:
@@
We can expect amplification if one of the $\lambda_{2,3}$ has modulus greater than unity. Denoting $\xi\mathrel{\mathop{:}}=c_{1}c_{2}(w_{x}^{2}+w_{y}^{2})$ we can write:
@@
Observe that the product of the two equals one:
$\lambda_{2}*\lambda_{3}=(1+0.5\xi)^{2}-0.5^{2}(\xi^{2}+4\xi)=1.$ (36)
Therefore, $\lambda_{2}=1/\lambda_{3}$. From this it follows that if $|\lambda_{2}|<1$ then $|\lambda_{3}|>1$ or vice-versa and an unstable mode has been found. We must only exclude the possiblity of $|\lambda_{2}|=|\lambda_{3}|=1$. However, in this case we can write $\lambda_{2}=e^{i\alpha},\lambda_{3}=e^{-i\alpha}$ and calculate:
@@
As emphasised earlier, this is by no means a formal proof of instability of any meshless approach, but instead merely an intuition as to where the locally unstable modes may come from. The analysis was limited to an interleaved leapfrog time-stepping, with no field staggering and a general expression of a derivative as a linear combination of stencil values. This setup includes our proposed RBF-FD and RBF-VFD based methods.